OpenStudy (anonymous):

The Hydro water department has a monthly service charge of $11.60 and a volume charge of $2.17 for every 100 cubic feet of water. The total monthly cost is equal to the service charge plus the charge for the volume used during the month. If we let x represent the number of cubic feet of water (in hundreds) and y represent the total monthly cost, then we have: total monthly cost = volume charge + service charge y = 2.17x + 11.60, If the Johnson family used 200 ft3 more water this month than last month, then they can expect

OpenStudy (anonymous):

a decrease in their water bill of $23.20. b a decrease in their water bill of $4.34. c an increase in their water bill of $4.34. d an increase in their water bill of $23.20.

OpenStudy (pfenn1):

We are given y=2.17x + 11.60 where x and y are given above. So if they use 200ft3 more water this month, the equation becomes y=2.17(x + 200) +11.60 So how does y change? Can you multiply it out?

OpenStudy (pfenn1):

Opps. I missed the fact that the 2.17 charge is per 100 cubic feet of water. y=2.17[x + (200/100)] + 11.60

OpenStudy (anonymous):

i still dont understand?

OpenStudy (pfenn1):

Do you understand the equation?

OpenStudy (anonymous):

nope

OpenStudy (pfenn1):

Okay. Just as it says above: The total cost, y, is equal to the monthly service charge, $11.60 (or the amount you are charged before you use any water) plus the volume charge of $2.17 per 100 cubic feet of water times the number of 100 cubic feet used, x. In equation form this is given by y=2.17x + 11.60

OpenStudy (anonymous):

a b c or d

OpenStudy (pfenn1):

But if x is the number of 100 cubic feet the family used this month and the family used 200 cubic feet more the next month, how many more 100 cubic feet did the family use the next month x + 200/100=x +2

OpenStudy (pfenn1):

I can tell you the answer but you need to understand how to figure it out. Substitute x + 2 into the first equation for x and you get y= 2.17 (x+2) + 11.60 = 2.17x + 4.34 + 11.60

OpenStudy (anonymous):

ok